OpenStudy (anonymous):

∫ (3x²+4-2)dx Can anyone walk me through this?

OpenStudy (anonymous):

the answer is 6x

OpenStudy (anonymous):

if u take the derivative of 4dx - 2dx u'll get nothing

OpenStudy (callisto):

∫ (3x²+4-2)dx = ∫ (3x²+2)dx = 3(1/3)x^3 +2(x) +c = x^3 +2x +c

OpenStudy (callisto):

i doubt if your question is ∫ (3x²+4x-2)dx ...

OpenStudy (anonymous):

gahh damnit i mistyped the problem. supposed to be +4x not +4. Sorry

OpenStudy (anonymous):

oh so its 6x+4

OpenStudy (anonymous):

\[\int (3x²+4-2)dx =\frac{3x^{2+1}}{2+1}+\frac{4x^{1+1}}{1+1}-2x^{0+1}\] \ \[=x^3+2x^2-2x\]

OpenStudy (anonymous):

∫ (3x²+4x-2)dx thats what the problem should look like

OpenStudy (callisto):

satellite73 helps you solve your problem :)

OpenStudy (anonymous):

Thanks for the help everyone